OpenStudy (anonymous):

3(x+1)^5-4(x+1)^6 using short cut

OpenStudy (anonymous):

(x+1)^5*(-4x-1)

OpenStudy (anonymous):

not times

OpenStudy (anonymous):

What exactly are you looking for? Because that is correct :D

OpenStudy (anonymous):

no its not right sory solve for 3(x+1)^5 and 4(x+1)^6 thats how are teacher showed us but i keep doing them wrong.

OpenStudy (anonymous):

We need an equation to solve for each :D

OpenStudy (anonymous):

equation ,you just have to sove it like a polynominal

OpenStudy (anonymous):

You mean expand it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OHH: 3(x+1)^5 = 3x^5 +15x^4+30x^3 + 30x^2 + 15x +3 4(x+1)^6 = 4x^6 + 24x^5 + 60x^4 + 80x^3 +60 x^2 + 24x =4 = -(x^6+21x^5+45x64+50x^3+30x^2+9x+1)